Solve (dy)/(dx)=(x-y+2)/(x-y-2)using an appropriate substitution.


1
Expert's answer
2021-07-20T17:28:47-0400

Solution

New variable u = x-y  =>   du/dx = 1-dy/dx  =>  du/dx = 1-(u+2)/(u-2)  => du/dx = -4/(u-2)

It is separable differential equation

(u-2)du = -4dx   =>  ∫(u-2)du = -4∫dx   =>  u2/2-2u = -4x+C  =>  u2-4u = -8x+C  =>  

"u(x)=2\pm\sqrt{4-8x+C}"

"y(x)=x-2\pm2\sqrt{1-2x+C}"